OpenStudy (anonymous):

Prove the identity: cos(2theta)/cos squared theta=1-tan squared theta

OpenStudy (anonymous):

\[\cos2x/\cos ^{2}x=1-\tan ^{2}x\] \[(\cos ^{2}x-\sin ^{2}x)/\cos ^{2}x=1-\tan ^{2}x\] \[\cos ^{2}x/\cos ^{2}x- \sin ^{2}x/\cos ^{2}x= 1-\tan ^{2}x\] \[1-\tan ^{2}x=1-\tan ^{2}x\]